use substitution
y=5x-7
-4x=+y=-1

That is the exact problem in my text book and to use the automated part of this website can you please tell me what to put in the slots

Answer by shree840(260)

You have two equations. we substitute y=5x-7 in the 2nd eq
-4x+y=-1( i believe this should be your eq. check your eq)
-4x+5x-7=-1
x-7=-1
x=6